Solve for x the following quadratic equations by factorising
X squared -6x+8=0

Answer by ewatrrr(24785)

x^2 - 6x + 8 = 0
(x-4)(x-2) = 0
(x-4)= 0, x = 4
(x-2) = 0 , x = 2
x = 2,4